OpenStudy (anonymous):

If f(x) = 2x 2 + 3, find f(3). A.31 B.15 C.21 D.39

OpenStudy (s3a):

You just replace all x variables with constants, 3.: f(3) = 2*(3)^2 + 3 = 2*9 + 3 = 18 + 3 = 21